How many mL of concentrated HCL solution are needed to provide 90.0g of HCL? Concentrated HCL solution is 37.0% HCl by weight and has a density of 1.19 g/mL.
1
Expert's answer
2017-09-07T09:43:07-0400
p =m/V
1,19 = 90/V
V = 75,63 ml - it's volume of 37% HCL
Need to solve a mass of full solution ( x = V(HCL)+(H2O) ), then
75,63 ml - 37%
x - 100%.
x = 75,63x100/37 = 204,4 ml (volume of 37% concentrated HCL).